OpenStudy (anonymous):

What is the slope-intercept form of the function described by this table? X: 1 2 3 4 Y: -2 -6 -10 -14 y=__x+___

OpenStudy (danjs):

y = m * x + b the first blank is the slope of the line m. You calculate that using two points from the table (x,y) and (x1,y1). The slope is found by..\[m = \frac{ x-x1 }{ y-y1 }\]

OpenStudy (danjs):

(x,y) = (1 , -2) (x,1 , y1) = (2, -6)

OpenStudy (danjs):

Calculate the slope from putting those numbers into the equation for slope(m). Then to form an equation for the line. Use the slope and one point from the line. (x1,y1). Point- slope form of a line: y - y1 = m*(x - x1)

OpenStudy (anonymous):

1-(-2) /2-(-6) = 3/8 = 0.375 im not very good with slope

OpenStudy (danjs):

Using the two points i put down above. you get \[slope ~m = \frac{ y-y1 }{ x-x1 } = \frac{ -2-(-6) }{ 1-2 } = \frac{ 4 }{ -1 } = -4\]

OpenStudy (danjs):

Now using that slope m=-4, and the point (x1,y1) = (2,-6) put that into the point slope form for a line y-y1 = m(x-x1) y - (-6) = -4*(x - 2) now rearrange to y=mx+b form y + 6 = -4x + 8 y = -4x + 8 - 6 y = -4x + 2